The minute hand of a clock is 5 cm long. Find the distance covered by its tip in 1 day take pi as 3.14
Answer by CharlesG2(834)

The minute hand of a clock is 5 cm long. Find the distance covered by its tip in 1 day take pi as 3.14
one end of the minute hand is at the center of a circle with radius 5 cm,
the tip of the minute hand touches the edge of this circle,
we need to find the circumference using 3.14 as our value for pi
circumference C = 2 * pi * radius r
C = 2 * 3.14 * 5
C = 10 * 3.14
C = 31.4 cm